OpenStudy (anonymous):

what is the number of real solutions: -11x^2 =x+11 (1 point)

ganeshie8 (ganeshie8):

put that in standard form and find the discriminant

OpenStudy (anonymous):

standard form is ax^2 + bx +c? how do i know which one is a b or c

OpenStudy (anonymous):

11x^2 + x + 11 = 0

OpenStudy (anonymous):

how do i find the discriminant?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

ganeshie8 (ganeshie8):

ax^2 + bx + c = 0 discrimant = b^2 - 4ac

OpenStudy (anonymous):

1^2 - 4(11x11) 1- 484 -483

ganeshie8 (ganeshie8):

yes, so the discriminant is negative what does that tell you about the solutions of a quadratic ?

OpenStudy (anonymous):

theres 2 solutions?

ganeshie8 (ganeshie8):

nope, google

OpenStudy (anonymous):

imaginary/ no real solution

ganeshie8 (ganeshie8):

yep!

OpenStudy (anonymous):

Thanks for your help :)